# Low-Overhead Channel Estimation via 3D Extrapolation for TDD mmWave Massive MIMO Systems Under High-Mobility Scenarios

### v1

@@ IV-B1 Spatial-Frequency Sampling Embedding
Due to the large number of antennas and subcarriers, the downlink channels’ spatial and frequency domain representation must be compressed to reduce the computational complexity of downlink channel extrapolation. Therefore, we propose the spatial-frequency sampling embedding layer. The spatial-frequency sampling embedding layer consists of one rearrangement operation, two LN operations, and one linear layer, as Fig. 4 (a) shows. The rearrangement operation splits the $N_{T}$ transmit antennas and $N_{c}$ subcarriers into $N_{1}$ transmit antenna groups and $N_{2}$ subcarrier groups, respectively. Each transmit antenna group has $\frac{N_{T}}{N_{1}}$ antennas and each subcarrier group has $\frac{N_{c}}{N_{2}}$ subcarriers. Note that since the number of UE antennas, i.e., $N_{R}$, is generally small, it is not necessary to further split UE antennas. Specifically, the rearrangement operation rearranges the sample dimension (known as the batch size $N_{b}$), $N_{1}$, and $N_{2}$ together to form a new sample dimension, and consequently reforms a representation dimension of length $2\times N_{R}\times\frac{N_{T}}{N_{1}}\times\frac{N_{c}}{N_{2}}$. Then, the dimension of the reformed representation is normalized by the first LN operation. Following that, the dimension is further projected to an embedding dimension $d_{emb}$ by the linear layer and normalized by the second LN operation. Straightforwardly, the spatial-frequency sampling embedding layer samples the spatial and frequency domain representation of downlink channels with a spatial sampling factor $N_{1}$ and a frequency sampling factor $N_{2}$ to reduce the representation dimension by resorting to the spatial and frequency correlations of downlink channels. Meanwhile, the $N_{1}$ sampled transmit antenna groups and $N_{2}$ sampled subcarrier groups are combined to facilitate model training. It is worth noting that after the generative Transformers, the inverse spatial-frequency sampling embedding layer is applied to transform the sampled representation dimension back to the original spatial and frequency domain representation. In the inverse spatial-frequency sampling embedding layer, all layers and operations are placed and operated in the inverse direction to the spatial-frequency sampling embedding layer, as Fig. 4 (b) shows.
@@ IV-B2 Generative Transformer and Masked MHSA Mechanism
The generative Transformer is one kind of autoregressive generative models that was proposed for text generation and is known as one of the most important driven forces for the transformative ChatGPT. The core of the generative Transformer is the masked MHSA mechanism, a pivotal component enabling the effective processing of input sequences. The masked MHSA mechanism allows the model to capture long-term dependencies among elements of input sequences and selectively attend to these elements while preventing it from looking ahead during training by masking future elements. To learn temporal dependencies efficiently for downlink channel extrapolation, we propose the generative Transformer-based downlink channel extrapolation network. The architecture shown in Fig. 4 (c) depicts that the proposed generative Transformer consists of a temporal positional encoding layer and $N_{G}$ generative Transformer layers. The temporal positional encoding layer is similar to (13) - (14). Each generative Transformer layer further consists of a masked MHSA sub-layer, a feed-forward (FF) sub-layer, and also RC and LN operations around each of the two sub-layers. Compared with the MHSA sub-layer, a masking operation is introduced before the query, key, and value projections in the masked MHSA sub-layer. The output of the masking operation, denoted by $\tilde{\mathbf{X}}^{M}\in\mathbb{R}^{N_{I}^{\prime}\times d_{R}^{\prime}}$ with $N_{I}^{\prime}$ being the number of input elements and $d_{R}^{\prime}$ being the representation dimension of each input element, can be expressed by
@@ V Simulations
In this section, we first introduce the simulation setup for numerical evaluations. Then, we compare the proposed KDD-SFCEN with traditional and DNN-based extrapolation methods under various spatial and frequency compression ratios, SNRs, and UE velocity settings. We then compare the TUDCEN with DNN-based temporal channel extrapolators. Finally, we present and analyze the overall performance of the proposed framework in terms of the achievable sum-rate of the system.
@@ V-A Simulation Setup
V-A1 Simulation Dataset and Parameters
The simulation dataset is constructed using the MATLAB 5G toolbox [29]. We consider a mmWave massive MIMO system with one BS and one single user. Uniform linear arrays (ULAs) are employed at the BS and the UE with $N_{T}=32$ and $N_{R}=4$. The system works in the TDD mode and operates with the OFDM modulation. The system comprises $52$ RBs, and each RB is formed by $12$ subcarriers and $14$ OFDM symbols according to the 5G specification [21]. In addition, the system follows the 5G frame structure shown in Fig. 1 (e) with $\mu=3$ and $N_{slot}=8$, and note that the SRS period $T_{p}=1$ ms (i.e., once per sub-frame). The SRS is configured at the last symbol index of each special slot. We adopt the clustered delay line (CDL)-B MIMO channel model to generate channel instances[30] at various UE mobility settings. $95$ channel instances are generated for constructing training and validation samples, and the BS transmits the SRS to the UE via these channel instances at the special slot of each sub-frame in $10$ frames (i.e., in total $100$ sub-frames / $800$ slots) at a signal-to-noise ratio (SNR) level of $5$ dB (i.e., $\gamma_{SNR}=5$ dB) and the UE mobility of 60 km/h (i.e., $v=60$ km/h). $5$ channel instances are generated for constructing testing samples, and the BS transmits the SRS to the UE via these channel instances at the special slots of $100$ sub-frames at different SNR levels. The SRS pattern is configured based on the spatial compression ratio $R_{s}$ and the frequency compression level $R_{f}$ as depicted in Section II-C. As a result, for a given ($R_{s}$, $R_{f}$) pair, $9,500$ samples are generated for training ($9,000$ samples) and validation ($500$ samples); for a given ($R_{s}$, $R_{f}$, $\gamma_{SNR}$, $v$) pair, $500$ samples are generated for testing. The spatial compression ratios, frequency compression ratios, SNR levels, UE mobility, and other key system parameters are listed in Table I. The hyper-parameter settings of the proposed KDD-SFCEN and the TUDCEN are shown in Table II and Table III, respectively.
@@ V-A2 Evaluation Metrics
where $N_{test}$ is the number of testing samples. Specifically, for uplink channel estimation, $\overline{\mathbf{H}}$ refers to $\mathbf{H}$ and $t=0$; for uplink-downlink channel calibration, $\overline{\mathbf{H}}$ refers to $\mathbf{H}^{d}$ and $t=1$; for downlink channel extrapolation, $\overline{\mathbf{H}}$ refers to $\mathbf{H}^{d}$, and $t=2,3,\ldots,N_{slot}-1$.
In addition, the achievable sum-rate in bps/Hz is adopted to evaluate the system performance, which is defined as
$R\{t\}=\sum_{j=1}^{N_{test}}\sum_{s=1}^{N_{sf}}\log_{2}\left(\left|1+\frac{\mathbf{H}^{d}\{s,t\}\mathbf{F}\{s,t\}(\mathbf{H}^{d}\{s,t\}\mathbf{F}\{s,t\})^{H}}{N_{R}\sigma_{n}^{2}}\right|\right),$ (54)
where $\sigma_{n}^{2}$ is the noise power.
@@ V-B Performance Evaluation of Proposed KDD-SFCEN
It can be seen from Fig. 6 and Fig. 6 that when $R_{f}=8$ or $R_{s}=4$ (corresponding to a $4$-fold reduction in pilot training overhead compared to $R_{f}=2$ or $R_{s}=1$), the NMSE performance of uplink channel estimation with the proposed KDD-SFCEN is around $-9$ dB and $-10$ dB, respectively, which is already good enough for uplink channel estimation. However, since the downlink CSI at future slots has to be extrapolated based on the estimated uplink CSI, it would be better to improve the performance of uplink channel estimation as much as possible. From Fig. 6 and Fig. 6, increasing $R_{f}$ from $2$ to $4$ or increasing $R_{s}$ from $1$ to $2$ leads to a minimum performance degradation compared to other increasing cases in $R_{f}$ or $R_{s}$. [^6] Note that from Fig. 6 that increasing $R_{f}$ leads to a near linear NMSE (in dB) performance degradation, thus the actual performance degradation increases in this progress. These actually reflect the marginal effects of pilots in either the frequency domain or the spatial domain. Therefore, it is promising that we combine $R_{f}=4$ and $R_{s}=2$ for joint spatial and frequency domain channel extrapolation to improve the uplink channel estimation performance and restrict the performance degradation due to the reduction of pilot symbols in the frequency domain and the spatial domain. Fig. 8 shows the NMSE performance of uplink channel estimation versus $\gamma_{SNR}$ at $R_{f}=4$ and $R_{s}=2$ (solid lines), where we further compare the proposed KDD-SFCEN with the SF-CNN[19] in addition to the linear spline interpolation method[9] and the DFT-interpolation method[9]. The simulation results show that the NMSE performance of the KDD-SFCEN reaches near $-12$ dB at $\gamma_{SNR}=20$ dB and the KDD-SFCEN outperforms the SF-CNN about $4$ dB and the traditional baselines about $9$ dB at $\gamma_{SNR}\geq 0$ dB. Even with a low SNR, i.e., $\gamma_{SNR}=-5$ dB, the KDD-SFCEN can achieve a performance enhancement around $3.5$ dB and $5$ dB compared with the SF-CNN and traditional baselines. Moreover, the NMSE performance of the linear spline interpolation method, the DFT-interpolation method, and SF-CNN at $R_{f}=2$ and $R_{s}=1$ are also shown in Fig. 8 with dotted lines, indicating that the KDD-SFCEN can reduce the pilot training overhead up to $4$ times compared to existing methods. These results demonstrate that combining spatial and frequency domain channel extrapolation can further improve the channel estimation performance ($2$ dB to $3$ dB) with the same pilot training overhead (i.e., a $4$-fold reduction in pilot training overhead).
@@
We further examine the robustness to the UE velocity of the proposed KDD-SFCEN. As depicted in Fig. 8, DNN-based methods (including the proposed KDD-SFCEN and the SF-CNN) achieve the best performance when the testing UE velocity is equal to the training UE velocity (i.e., $v=60$ km/h), showing their capability in capturing spatial-frequency domain features and improving channel estimation accuracy. When the testing UE velocity decreases or increases, the performance of the DNN-based methods degrades. Moreover, the greater the deviation of the testing UE velocity from the training UE velocity, the more significant the performance degradation will be. Nonetheless, the proposed KDD-SFCEN achieves sufficient channel estimation accuracy (around or above $-8$ dB) and outperforms all baselines with $v\leq 90$ km/h at $R_{f}=4$ and $R_{s}=2$, showing its outstanding robustness to the UE velocity.
@@ V-C Performance Evaluation of Proposed TUDCEN
Figure 11: The achievable sum-rate performance of slot-level channel extrapolation versus the index of a slot in a sub-frame at $R_{f}=4$, $R_{s}=2$, $\gamma_{SNR}=20$ dB, and $v=60$ km/h.
We then evaluate the performance of slot-level channel extrapolation with the proposed TUDCEN and a baseline temporal channel extrapolation method. i.e., the LSTM-based channel predictor[17]. Note that the uplink-downlink channel calibration is applied to all methods. We further combine the proposed KDD-SFCEN and the baseline uplink channel estimation method SF-CNN with the proposed TUDCEN and the LSTM. The NMSE performance and the achievable sum-rate performance at $R_{f}=4$, $R_{s}=2$, $\gamma_{SNR}=20$ dB, and $v=60$ km/h from the first downlink slot to the seventh slot are shown in Fig. 11 and Fig. 11, respectively.
It can be seen from Fig. 11 that without slot-level channel extrapolation (dotted lines), there are significant estimation errors between the estimated downlink channel at the first downlink slot and the downlink channels at other downlink slots. Therefore, slot-level channel extrapolation is essential to reduce the channel estimation errors for these downlink channels. From Fig. 11, the ‘KDD-SFCEN + TUDCEN’ (the orange solid line) always achieves the best NMSE performance compared with other methods. In addition, the ‘SF-CNN + TUDCEN’ (the azure solid line) also keeps a relatively good extrapolation accuracy, showing that the proposed TUDCEN can also be applied to other uplink channel estimation methods. Compared with the proposed TUDCEN, the performance of the ‘KDD-SFCEN + LSTM’ and ‘SF-CNN + LSTM’ (dashed lines) degrades sharply to a poor NMSE regime at the second slot, indicating that the LSTM-based channel predictor is unable to deal with high-dimensional CSI matrices and provide accurate slot-level channel estimates.
As depicted in Fig. 11, both the ‘KDD-SFCEN + TUDCEN’ and ‘SF-CNN + TUDCEN’ (orange and azure solid lines) achieve a good achievable sum-rate performance at all downlink slots. Specifically, the ‘KDD-SFCEN + TUDCEN’ achieves $89\%$ (at the seventh downlink slot) to $99\%$ (at the first downlink slot) of the upper bound achievable sum-rate obtained with the perfect CSI (the red solid line). It is also worth emphasizing that the worst achievable sum-rate performance obtained by the ‘KDD-SFCEN + TUDCEN’ is near the achievable sum-rate performance that can be obtained by traditional linear spline interpolation and DFT interpolation methods with a small SRS period $T_{p}=0.25$ ms (purple and yellow solid lines), demonstrating that with the proposed framework, the pilot signal can be sent less frequently (i.e., $T_{p}=1$ ms) while maintaining a satisfying achievable sum-rate performance, thereby further reducing the pilot training overhead by $4$ times. In addition, compared with the general pilot training scheme (Fig. 1 (d)), the proposed TUDCEN-based slot-level channel extrapolation-aided pilot training scheme (Fig. 1 (e)) configures more slots for downlink data transmission instead of for uplink channel estimation, thereby significantly improving the system’s spectral efficiency.

### v2

@@ IV-B1 Spatial-Frequency Sampling Embedding
Due to the large number of antennas and subcarriers, the downlink channels’ spatial and frequency domain representation must be compressed to reduce the computational complexity of downlink channel extrapolation. Therefore, resorting to spatial and frequency correlations of wireless channels, we propose the spatial-frequency sampling embedding layer that can reduce the spatial and frequency dimensions of high-dimensional CSI matrices, thereby reducing the computational complexity. The spatial-frequency sampling embedding layer consists of one rearrangement operation, two LN operations, and one linear layer, as Fig. 4 (a) shows. The rearrangement operation splits the $N_{T}$ transmit antennas and $N_{c}$ subcarriers into $N_{1}$ transmit antenna groups and $N_{2}$ subcarrier groups, respectively. Each transmit antenna group has $\frac{N_{T}}{N_{1}}$ antennas and each subcarrier group has $\frac{N_{c}}{N_{2}}$ subcarriers. Note that since the number of UE antennas, i.e., $N_{R}$, is generally small, it is not necessary to further split UE antennas. Specifically, the rearrangement operation rearranges the sample dimension (known as the batch size $N_{b}$), $N_{1}$, and $N_{2}$ together to form a new sample dimension, and consequently reforms a representation dimension of length $2\times N_{R}\times\frac{N_{T}}{N_{1}}\times\frac{N_{c}}{N_{2}}$. Then, the dimension of the reformed representation is normalized by the first LN operation. Following that, the dimension is further projected to an embedding dimension $d_{emb}$ by the linear layer and normalized by the second LN operation. Straightforwardly, the spatial-frequency sampling embedding layer samples the spatial and frequency domain representation of downlink channels with a spatial sampling factor $N_{1}$ and a frequency sampling factor $N_{2}$ to reduce the representation dimension by resorting to the spatial and frequency correlations of downlink channels. Meanwhile, the $N_{1}$ sampled transmit antenna groups and $N_{2}$ sampled subcarrier groups are combined to facilitate model training. It is worth noting that after the generative Transformers, the inverse spatial-frequency sampling embedding layer is applied to transform the sampled representation dimension back to the original spatial and frequency domain representation. In the inverse spatial-frequency sampling embedding layer, all layers and operations are placed and operated in the inverse direction to the spatial-frequency sampling embedding layer, as Fig. 4 (b) shows.
@@ IV-B2 Generative Transformer and Masked MHSA Mechanism
The generative Transformer is one kind of autoregressive generative models that was proposed for text generation and is known as one of the most important driven forces for the transformative ChatGPT. The core of the generative Transformer is the masked MHSA mechanism, a pivotal component enabling the effective processing of input sequences. The masked MHSA mechanism allows the model to capture long-term dependencies among elements of input sequences and selectively attends to these elements while preventing it from looking ahead during training by masking future elements. To learn temporal dependencies efficiently for downlink channel extrapolation, we propose the generative Transformer-based downlink channel extrapolation network. The architecture shown in Fig. 4 (c) depicts that the proposed generative Transformer consists of a temporal positional encoding layer and $N_{G}$ generative Transformer layers. The temporal positional encoding layer is similar to (13) - (14). Each generative Transformer layer further consists of a masked MHSA sub-layer, a feed-forward (FF) sub-layer, and also RC and LN operations around each of the two sub-layers. Compared with the MHSA sub-layer, a masking operation is introduced before the query, key, and value projections in the masked MHSA sub-layer. The output of the masking operation, denoted by $\tilde{\mathbf{X}}^{M}\in\mathbb{R}^{N_{I}^{\prime}\times d_{R}^{\prime}}$ with $N_{I}^{\prime}$ being the number of input elements and $d_{R}^{\prime}$ being the representation dimension of each input element, can be expressed by
@@ V Simulations
In this section, we first introduce the simulation setup for numerical evaluations. Then, we compare the proposed KDD-SFCEN with traditional and DNN-based extrapolation methods under various spatial and frequency compression ratios, SNRs, and UE velocity settings. We then compare the TUDCEN with DNN-based temporal channel extrapolators. Finally, we present and analyze the overall performance of the proposed framework in terms of the achievable rate of the system.
@@ V-A Simulation Setup
V-A1 Simulation Dataset, Parameters, and Model Training
The simulation dataset is constructed using the MATLAB 5G toolbox [28]. We consider a mmWave massive MIMO system with one BS and one single user. Uniform linear arrays (ULAs) are employed at the BS and the UE with $N_{T}=32$ and $N_{R}=4$. The system works in the TDD mode and operates with the OFDM modulation. The system comprises $52$ RBs, and each RB is formed by $12$ subcarriers and $14$ OFDM symbols according to the 5G specification [22]. In addition, the system follows the 5G frame structure shown in Fig. 1 (e) with $\mu=3$ and $N_{slot}=8$, and note that the SRS period $T_{p}=1$ ms (i.e., once per sub-frame). The SRS is configured at the last symbol index of each special slot. We adopt the clustered delay line (CDL)-B MIMO channel model to generate channel instances[29] at various UE mobility settings. The CDL-B channel model uses multiple clusters to characterize multipath propagation and supports a frequency range from $0.5$ GHz to $100$ GHz, which is widely adopted for link-level evaluation. $95$ channel instances are generated for constructing training and validation samples, and the BS transmits the SRS to the UE via these channel instances at the special slot of each sub-frame in $10$ frames (i.e., in total $100$ sub-frames / $800$ slots) at a signal-to-noise ratio (SNR) level of $5$ dB (i.e., $\gamma_{SNR}=5$ dB) and the UE mobility of 60 km/h (i.e., $v=60$ km/h). $5$ channel instances are generated for constructing testing samples, and the BS transmits the SRS to the UE via these channel instances at the special slots of $100$ sub-frames at different SNR levels. The SRS pattern is configured based on the spatial compression ratio $R_{s}$ and the frequency compression level $R_{f}$ as depicted in Section II-C. As a result, for a given ($R_{s}$, $R_{f}$) pair, $9,500$ samples are generated for training ($9,000$ samples) and validation ($500$ samples); for a given ($R_{s}$, $R_{f}$, $\gamma_{SNR}$, $v$) pair, $500$ samples are generated for testing. The spatial compression ratios, frequency compression ratios, SNR levels, UE mobility, and other key system parameters are listed in Table I. The hyper-parameter settings of the proposed KDD-SFCEN and the TUDCEN are shown in Table II and Table III, respectively. The model training processes are as follows. The KDD-SFCEN for uplink channel estimation is first trained with $L_{1}$ in (37), followed by training the UDCCN for uplink-downlink calibration with $L_{2}$ in (51), and finally, the DCEN for downlink channel extrapolation is trained with $L_{3}$ in (52), all based on the supervised learning paradigm. Specifically, $\mathbf{H}{\{s,0\}}$, ${\mathbf{H}^{d}}\{s,1\}$, and $\mathbf{H}^{d}\{s,t\}$ for $t=2,3,\ldots,N_{slot}-1$ are provided as labels for training the three models, respectively. By minimizing the three loss functions on the training samples with the Adam optimizer [30] and validating the models on the validation samples, all three models are effectively trained and validated.
@@ V-A2 Evaluation Metrics
where $N_{test}$ is the number of testing samples, and $\{j,s,t\}$ refers to the $t$-th slot of the $s$-th sub-frame of the $j$-th testing sample. Specifically, for uplink channel estimation, $\overline{\mathbf{H}}$ refers to $\mathbf{H}$ and $t=0$; for uplink-downlink channel calibration, $\overline{\mathbf{H}}$ refers to $\mathbf{H}^{d}$ and $t=1$; for downlink channel extrapolation, $\overline{\mathbf{H}}$ refers to $\mathbf{H}^{d}$, and $t=2,3,\ldots,N_{slot}-1$.
In addition, the achievable rate in bps/Hz is adopted to evaluate the system performance, which is defined as (54) at the top of the next page, where $\sigma_{n}^{2}=0.01$ is the noise power, and $\mathbf{F}^{RF}_{i}\in\mathbb{C}^{N_{T}\times N_{RF}}$ and $\mathbf{F}^{BB}_{i}\in\mathbb{C}^{N_{RF}\times N_{R}}$ are the analog precoding matrix and digital precoding matrix for the $i$-th subcarrier designed based on the estimated downlink CSI $\hat{\mathbf{H}}^{d}_{i}$, respectively. [^6] Note that since the analog precoder can only control the phase of the signals and are typically frequency-independent, the same analog precoder will be applied to all subcarriers in practice when designing $\mathbf{F}^{RF}_{i}$.
$R\{t\}=\frac{1}{N_{test}N_{sf}N_{c}}\sum_{j=1}^{N_{test}}\sum_{s=1}^{N_{sf}}\sum_{i=1}^{N_{c}}\log_{2}\left(\left|\mathbf{I}+\frac{\mathbf{H}^{d}_{i}\{j,s,t\}\mathbf{F}^{RF}_{i}\{j,s,t\}\mathbf{F}^{BB}_{i}\{j,s,t\}(\mathbf{F}^{BB}_{i}\{j,s,t\})^{H}(\mathbf{F}^{RF}_{i}\{j,s,t\})^{H}(\mathbf{H}^{d}_{i}\{j,s,t\})^{H}}{N_{R}\sigma_{n}^{2}}\right|\right).$ (54)
@@ V-B Performance Evaluation of Proposed KDD-SFCEN
It can be seen from Fig. 5 and Fig. 6 that when $R_{f}=8$ / $R_{s}=4$ (corresponding to a $4$-fold reduction in pilot training overhead compared to the minimum frequency / spatial compression ratio $R_{f}=2$ / $R_{s}=1$ according to Table I), the NMSE performance of uplink channel estimation with the proposed KDD-SFCEN is around $-9$ dB / $-10$ dB at $\gamma_{SNR}=20$ dB, which is already good enough for uplink channel estimation. However, since the downlink CSI at future slots has to be extrapolated based on the estimated uplink CSI, it would be better to improve the performance of uplink channel estimation as much as possible. From Fig. 5 and Fig. 6, increasing $R_{f}$ from $2$ to $4$ or increasing $R_{s}$ from $1$ to $2$ leads to a minimum performance degradation compared to other increasing cases in $R_{f}$ or $R_{s}$. [^7] Note that from Fig. 5 that increasing $R_{f}$ leads to a near linear NMSE (in dB) performance degradation, thus the actual performance degradation increases in this progress. These actually reflect the marginal effects of pilots in either the frequency domain or the spatial domain. Therefore, it is promising that we combine $R_{f}=4$ and $R_{s}=2$ for joint spatial and frequency domain channel extrapolation to improve the uplink channel estimation performance and restrict the performance degradation due to the reduction of pilot symbols in the frequency domain and the spatial domain. Fig. 7 shows the NMSE performance of uplink channel estimation versus $\gamma_{SNR}$ at $R_{f}=4$ and $R_{s}=2$ (solid lines), which can also reduce $4$ times of pilot training overhead. The proposed KDD-SFCEN obtains an NMSE of $-12$ dB at $\gamma_{SNR}=20$ dB, indicating that joint spatial-frequency domain channel extrapolation can obtain a $3$ / $2$ dB of performance gain compared with individual frequency / spatial domain channel extrapolation. In addition, we further compare the proposed KDD-SFCEN with the DRCN[31], the SF-CNN[19], the linear spline interpolation method[9], and the DFT-interpolation method[9] when $R_{f}=4$ and $R_{s}=2$ (solid lines) / $R_{f}=2$ and $R_{s}=1$ (dotted lines). The simulation results show that when $R_{f}=4$ and $R_{s}=2$, the KDD-SFCEN outperforms the DNN-based baselines (i.e., the DRCN and the SF-CNN) around $4$ dB and the traditional baselines around $9$ dB at $\gamma_{SNR}\geq 0$ dB; even with a low SNR, i.e., $\gamma_{SNR}=-5$ dB, the KDD-SFCEN can achieve a performance enhancement more than $3$ dB and $5$ dB compared with the DNN-based and traditional baselines, respectively. When $R_{f}=2$ and $R_{s}=1$, the KDD-SFCEN outperforms the DNN-based and traditional baselines more than $5$ dB and $11$ dB respectively at the high SNR regime, and about $3$ dB and $5$ dB respectively at the low SNR regime. Particularly, it can be found from Fig. 7 that the NMSE performance of the proposed KDD-SFCEN with $R_{f}=4$ and $R_{s}=2$ is even better than existing methods with $R_{f}=2$ and $R_{s}=1$, indicating that the KDD-SFCEN can reduce the pilot training overhead up to $4$ times compared to existing methods.
@@
We further examine the robustness of the proposed KDD-SFCEN to the UE velocity. As depicted in Fig. 8, DNN-based methods (including the proposed KDD-SFCEN, the DRCN, and the SF-CNN) achieve the best performance when the testing UE velocity is equal to the training UE velocity (i.e., $v=60$ km/h), showing their capability in capturing spatial-frequency domain features and improving channel estimation accuracy. When the testing UE velocity decreases or increases, the performance of the DNN-based methods degrades. Moreover, the greater the deviation of the testing UE velocity from the training UE velocity, the more significant the performance degradation will be. Nonetheless, the proposed KDD-SFCEN achieves sufficient channel estimation accuracy ($\leq-8$ dB) and outperforms all baselines with $v\leq 90$ km/h at $R_{f}=4$ and $R_{s}=2$, showing its outstanding robustness to the UE velocity.
@@ V-C Performance Evaluation of Proposed TUDCEN
Figure 11: The achievable rate performance of slot-level channel extrapolation versus the index of a slot in a sub-frame at $R_{f}=4$, $R_{s}=2$, $\gamma_{SNR}=20$ dB, and $v=60$ km/h.
We then evaluate the performance of slot-level channel extrapolation with the proposed TUDCEN and two temporal channel extrapolation methods, i.e., the TPreNet[32] and the LSTM-based channel predictor[17]. Note that the uplink-downlink channel calibration is applied to all methods. The NMSE performance and the achievable rate performance at $R_{f}=4$, $R_{s}=2$, $\gamma_{SNR}=20$ dB, and $v=60$ km/h from the first downlink slot to the seventh slot are shown in Fig. 10 and Fig. 11, respectively. For better comparison, the upper bound achievable rate obtained with the perfect CSI and the upper bound slot-average achievable rate with a small SRS period $T_{p}=0.25$ ms (i.e., only $4$ slots are available for downlink transmission) are also shown in Fig. 11.
It can be seen from Fig. 10 that without slot-level channel extrapolation (dotted lines), there are significant estimation errors for downlink channels at slots following the first downlink slot. Therefore, slot-level channel extrapolation is essential to reduce the channel estimation errors for these downlink channels. From Fig. 10 and Fig. 11, compared with the proposed ‘KDD-SFCEN + TUDCEN’, the performance of the ‘KDD-SFCEN + TPreNet’ and ‘KDD-SFCEN + LSTM’, in terms of either NMSE or achievable rate, degrades sharply to a poor regime at the second slot, indicating that these channel predictors are unable to deal with high-dimensional CSI matrices and provide accurate slot-level channel estimates. As depicted in Fig. 11, the proposed ‘KDD-SFCEN + TUDCEN’ can achieve near $80\%$ of the upper bound achievable rate even after $5$ slots from channel estimation. In addition, the achievable rate performance of the proposed ‘KDD-SFCEN + TUDCEN’ is always better than the upper bound slot-average achievable rate with a small SRS period $T_{p}=0.25$ ms, demonstrating that with the proposed 3D channel extrapolation framework, the pilot signal can be sent less frequently (i.e., $T_{p}=1$ ms) while maintaining a satisfying achievable rate performance, thereby further reducing the pilot training overhead by $4$ times.